SOLVE THE GIVEN DIFFERENTIAL EQUATION BY SEPARATION OF VARIABLES
DY/DX=Y-2/X-1
Y(2)=2
1
Expert's answer
2020-12-15T02:51:01-0500

Here the problem is written in lower case.

"\frac{dy}{dx}= \frac{y-2}{x-1}"

=> "\frac{dy}{y-2}= \frac{dx}{x-1}"

Integrating we get

"\int\frac{dy}{y-2}=\int \frac{dx}{x-1}"

ln|y-2| = ln|x-1| + ln|C|

=> ln|y-2| = ln|C(x-1)|

=>| y - 2 |= |C(x-1)|


By initial condition , y = 2 when x = 2

0 = |C| i.e C = 0

Therefore the particular solution is

|y - 2| = 0

i.e. y= 2